write a quadratic equation in the variable x having the given numbers as solutions. type the equation in standard form ax^2+bx+c=0. solutions are -5,2
Thank you!

Answer by oberobic(2304)

Given the solutions (zeroes) are -5 and 2. That means the quadratic equation is:
a(x+5)(x-2) = 0
.
We cannot say what the value of 'a' is unless we have some additional information.
'a' is simply a constant that would affect the graph's position.
.
(x+5)(x-2) = x^2 +3x -10
.
In the following graph:
The red line assumes 'a' = 1.
The green line assumes 'a' = -1.
The blue line assumes 'a' = 1/2.
The purple line assumes 'a' = -1/4.
.